X Company is considering buying a part in 2020 that it has been making for the past several years. A company has offered to supply this part for $15.39 per unit. Budgeted production in 2020 is 54,000 units, and budgeted per-unit production costs are: Materials $6.00 Direct labor [all variable] 4.30 Total overhead 4.60 Total $14.90 $75,600 of X Company's total overhead costs are fixed; $58,968 of the $75,600 are unavoidable even if it buys the part. Also, if X Company buys the part, it can rent the equipment that was used to make the part to another company for $75,000. 3. If X Company continues to make the part instead of buying it, it will save 1101282 Incorrect. Tries 2/3 Previous Tries 4. At what production level would X company be indifferent between making and buying the part?
Answer #1

1) Differential analysis

Make Buy
Direct material 324000
Direct labor 232200
Variable overhead 172800
Fixed overhead 16632
Opportunity Cost 75000
Purchase cost 831060
Total relevant cost 820632 831060

It will save = 831060-820632 = 10425

2) Indifferent point

Manufacturing cost = Purchase cost

13.50X+91632 = 15.39X

-1.89X = 91632

X(indifferent point) = 48483 Unit